# LLM-Generated Black-box Explanations Can Be Adversarially Helpful

### v1

@@ Abstract
Large Language Models (LLMs) are becoming vital tools that help us solve and understand complex problems by acting as digital assistants. LLMs can generate convincing explanations, even when only given the inputs and outputs of these problems, i.e., in a “black-box” approach. However, our research uncovers a hidden risk tied to this approach, which we call adversarial helpfulness. This happens when an LLM’s explanations make a wrong answer look right, potentially leading people to trust incorrect solutions. In this paper, we show that this issue affects not just humans, but also LLM evaluators. Digging deeper, we identify and examine key persuasive strategies employed by LLMs. Our findings reveal that these models employ strategies such as reframing the questions, expressing an elevated level of confidence, and cherry-picking evidence to paint misleading answers in a credible light. To examine if LLMs are able to navigate complex-structured knowledge when generating adversarially helpful explanations, we create a special task based on navigating through graphs. Some LLMs are not able to find alternative paths along simple graphs, indicating that their misleading explanations aren’t produced by only logical deductions using complex knowledge. These findings shed light on the limitations of black-box explanation setting. We provide some advice on how to use LLMs as explainers safely.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong explaining capabilities. They can generate texts containing logical steps towards solving complex problems (Trinh et al., 2024; Sprague et al., 2023), and incorporate users’ situations (Mondal et al., 2024; Zhu et al., 2023; Zhou et al., 2024). The literature in NLP has many works showing LLMs can generate explanations that are convincing (Wiegreffe et al., 2022), informative (Chen et al., 2023) and faithful (Lyu et al., 2023) to levels that are comparable to human-written explanations. These abilities lead to a promising use case: LLMs as assistants that explain problems to us.
Sounds like we have a convenient use case: we pass a complex problem and its answer to an LLM, together with a suitably-formatted prompt. Hopefully, this “black-box” approach will let the LLM generate an explanation that is instructive and helpful for us to understand the problem. However, as we will show in this paper, this approach may backfire. Specifically, when we prompt an LLM to explain an incorrect answer to a problem, this “black-box” approach may produce explanations that lead us to believe in the incorrect answers.
We first try to gauge the extent of this adversarially helpfulness problem. We let human annotators rate the extent they are convinced of an incorrect answer, before and after seeing an LLM-generated explanation. On a commonsense reasoning task, their ratings are significantly improved by explanations generated by three commercial-tier LLMs. We also use some proxy evaluator models to repeat the survey, and observe similar effects.
@@
Figure 1: Examples of LLM-generated explanations towards incorrect labels in a commonsense QA question (above) and an NLI question (below). We mark the persuasion strategies with color-codeed angle brackets. To save space, the parts that do not contain persuasion strategies are omitted with […].
Then, we proceed with studying the strategies of the adversarially helpful explanations. We identify ten strategies that we consider relevant to the explanations being adversarially helpful, and examine the strategies used by LLMs in generating these explanations. We find some alarmingly high frequencies. For example, over 90% of the explanations in inference problems involve reframing to varying extents, and over 60% of the explanations in the commonsense problems involve the selective presentation of either the facts or the evidence.
Then we dive into the mechanisms. Are these LLMs able to generate these adversarially helpful explanations because they are able to navigate through complex knowledge, like finding an alternate path in a graph? We set up a symbolic reasoning task, where we ask the LLMs to figure out an alternate path that leads to a specified destination node. This is an abstraction for “explaining an incorrect answer”. We find that the weaker models are unable to complete this task, especially when the graph complexity increases. These indicate that the generation of adversarially helpfulness explanations may involve more than the abilities in deductive reasoning and logical inference, which matches the prior observation that additional strategies (e.g., reframing) are used.
The findings shed light on the limitations of the black-box explanation setting, and motivates us to provide some recommendations for safer usage of LLMs as explaining assistants.
@@ 2 Related Works
Recently, a number of works have leveraged the reasoning abilities of LLMs: chain-of-thought (Nye et al., 2021; Wei et al., 2022), tree-of-thought (Yao et al., 2024), graph-of-thought (Besta et al., 2023) and everything-of-thought (Ding et al., 2023) are some examples. We follow their approaches to elicit the reasoning abilities from LLMs.
@@
Model-generated explanations can have significant impacts on both human users, for example in mitigating misinformation (Hsu et al., 2023), rescale human judgments (Wadhwa et al., 2023) and understand models’ behavior (Hase & Bansal, 2020; Chen et al., 2022). Researchers have tried to use explanations for model development, yet there are mixed results (Saha et al., 2024; Im et al., 2023).
@@
The explainability research of AI have identified pitfalls (those unintended effects) versus dark patterns (those resulting from intended misuse) (Ehsan & Riedl, 2021). The adversarial helpfulness of LLM-generated explanation is an example of explainability pitfalls. Specific to the explanations in natural language, these explanations are selective, and can be subjective or misleading (Kunz & Kuhlmann, 2024) or unreliable (Ye & Durrett, 2022). Adversarial helpfulness is different from other problems, in that it might sometimes be a feature; we elaborate in the discussion section.
@@
LLMs with self explanations can be biased towards incorrect answers by superficial patterns like the ordering of choices (Turpin et al., 2024). Our paper considers a different setting: instead of planting superficial patterns in the demonstrations, we instruct the LLMs to explain an incorrect answer in a zero-shot manner, resembling how a user would use the LLM as an explaining assistant. This paper is also relevant to the jailbreaking/red-teaming of LLMs (Zou et al., 2023; Zeng et al., 2024; Deng et al., 2023; Ganguli et al., 2022). Instead of developing the jailbreaking or defending algorithms, we focus on the problem itself, and study the strategies adopted by the LLMs when generating adversarially helpful explanations.
@@ 3 Experiment setup
Two datasets are used: ECQA (Aggarwal et al., 2021a) and SNLI (Bowman et al., 2015). For ECQA, we sampled 500 problems and hand-pick a “second-best answer” that we consider to be only slightly worse than the correct answer designated by the dataset. Anecdotally, we found that those more abstract problems lead to more significant “adversarial helpfulness” explanations, so we skipped some very concrete problems. For SNLI, we sampled 300 problems from each of the datasets with label Entailment and Contradictory, respectively. We found it was almost impossible to write sufficiently logical arguements (with humans or LLMs) that an Entailment sentence pair should be Contradictory (or vice versa), so we only considered the cases that argue for a Neutral label.
@@
We consider the following models as evaluators: Vicuna-33B-v1.3 (Chiang et al., 2023), WizardLM-70B-V1.0 (Xu et al., 2023), Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). These are the top-ranking [^1] As of February 2024. Note that Qwen1.5-72B-Chat (Bai et al., 2023) and Yi-34B-Chat (AI et al., 2024) are also ranked on the top places, but we only included one LLM per model size to reduce the computation cost. models on the lmsys Chatbot Arena leaderboard (Zheng et al., 2023) that provide weight access. The weight access is important, since it allows us to query the probabilities in setting up the automatic evaluation — we will describe the setup in Section 5.
@@
The “adversarially helpful” explanations are generated by prompting the LLM explainer to explain an incorrect answer to a question. Figure 2 lists the prompting templates we used. Empirically, we found other prompting templates can lead to explanations demonstrating a similar “adversarially helpful” behavior, as long as we specifically instruct the model to convince the audience of the incorrect answer.
@@ 5 Automatic evaluation of adversarial helpfulness
Note that the utility of LLM that evaluate the convincingness should be taken with caution. First, LLMs have demonstrated some evidence to model humans’ thoughts (i.e., “theory-of-mind” modeling) (Kosinski, 2024), but the actual capability is being debated. Second, in some cases the LLM-given ratings show some “degeneration” trends. For example, in NLI $E\rightarrow N$, all C_after scores computed by Mixtral-8x7B are the same, resulting in a average score of 3.00. This indicates that some of the scores given by the proxy evaluators might result from the dataset artefacts rather than the contents.
@@
When getting down to the item-wise level, the evaluators show some different trends. First, humans correlate weakly with the proxy models. On the dataset where human annotator results are available, we compute the Pearson correlations between the averaged human results and the evaluator models. None of the correlations are significant, indicating that the evaluator models show very different fluency, correctness and convincingness ratings from the human annotators. Second, human results are random. This is because the MTurk distribute the annotation tasks to more than three annotators, so the inter-annotator agreements do not apply here. Third, the proxy models assign different scores. We compute the Cohen’s Kappa between each pair of the proxy models. None of them have a value larger than 1 for any score, indicating poor agreements between the evaluators. This is because the proxy models have different “baseline marking guidelines”, as is illustrated by the drastically different mean scores on Table 1.
@@ 6 Strategies toward adversarial helpfulness
The recent literature involve many taxonomies of persuasion strategies. For example, Dimitrov et al. (2021) identified strategies in social media texts. Piskorski et al. (2023) considered news. Zeng et al. (2024) considered 40 techniques for jailbreaking LLMs as well as other NLP tasks. Inspired by these works, we identify ten persuasion strategies that we consider particularly relevant to LLM-generated explanations. A brief summarization of each strategy is included in Appendix B.
We use GPT-4 Turbo (gpt-4-0125-preview) to detect the persuasion strategies used in the explanations. Figure 1 lists some examples of the identified strategies.
@@
Table 2 lists the frequencies of the persuasion strategies adopted by the three explainer models. Some trends emerge from the results.
First, the sheer numbers of frequency are alarming. For the commonsense questions, more than $70\%$ of the explanations highlight the comparative advantage towards the incorrect answers. For inference tasks, more than $90\%$ of the explanations involve reframing of the questions. While these results are only on the commonsense and inference datasets, we expect that the real-world applications of natural language explanations frequently include these persuasion strategies.
Second, the three explainer models have common trends in applying persuasion strategies. The LLM-generated explanations appear overly confident. [^2] This may related to the overly-confident tone in our prompts, as listed in Figure 2. Strategies like selective evidence, and selective fact presentation are frequently used, and the strategies like appeal to authority and analogical evidence are very infrequently used in any of the models. These indicate that the safeguarding against the adversarially helpful explanations can benefit from the safeguarding against a finite set of persuasion strategies.
@@ 10. Complex inference 4 1 4 9 8 5 3 7 3
Table 2: The frequencies (out of 100) of persuasion strategies adopted by explainer models. The three mostly adopted strategies per column are marked in bold font.
@@ 7 A structural analysis towards adversarial helpfulness
We consider the process of explanation to be an instance of path finding on a graph. In this path finding problem, we find a path from the “root” node to the “answer” node. The explanation serves as the path that connects the known (the “root” node) to the unknown (the “answer” node).
@@
LLMs have limited capabilities in capturing the structural information in graphs (Huang et al., 2023a). Based on the intuition, we attach some simplification assumptions. First, all branches have the same path length. Second, each non-root node uniquely appears on only one path. Figure 4 shows an example of such graph with $B=3$ branches of $L=4$ path.
@@ 8 Discussion
There are intricate differences between these two improvement areas of LLM-generated explanations. Adversarial helpfulness refers to the explainer’s behavior that rationalizes an incorrect phenomenon. Unfaithfulness, on the other hand, rationalizes problems using the incorrect reasons or irrelevant facts. Using the analogy in the graph reasoning setting, an adversarially helpful explanation is analogous to a path (with or without an incorrect node) that leads to an incorrect destination, and an unfaithful explanation is analogous to a path that involves incorrect nodes or edges, yet leads to the correct destination.
@@
Plausibility is different from faithfulness — for an extended discussion of their difference please refer to Agarwal et al. (2024). Here we consider that adversarial helpfulness has an overlap with plausibility; the latter indicates a potential negative impact towards human, whereas the former does not. Plausibility is a feature, but adversarial helpfulness is a bug.
@@
We provide some recommendations here.
First, let humans make decisions, not AIs. We can use the AI explainer as a “Prudence” model (Miller, 2023). A “Prudence” model provides evidence supporting human’s decisions, without directly giving us an answer. The alternative, “Bluster” model, recommends an answer, optionally with the rationales supporting that answer. Since the rationales supporting the answer can be adversarially helpful, it is important to delegate the least possible amount of “decision-making” responsibility to the explainers (especially those based on generative AI).

### v2

@@ Abstract
Large Language Models (LLMs) are becoming vital tools that help us solve and understand complex problems by acting as digital assistants. LLMs can generate convincing explanations, even when only given the inputs and outputs of these problems, i.e., in a “black-box” approach. However, our research uncovers a hidden risk tied to this approach, which we call adversarial helpfulness. This happens when an LLM’s explanations make a wrong answer look right, potentially leading people to trust incorrect solutions. In this paper, we show that this issue affects not just humans, but also LLM evaluators. Digging deeper, we identify and examine key persuasive strategies employed by LLMs. Our findings reveal that these models employ strategies such as reframing the questions, expressing an elevated level of confidence, and cherry-picking evidence to paint misleading answers in a credible light. To examine if LLMs are able to navigate complex-structured knowledge when generating adversarially helpful explanations, we create a special task based on navigating through graphs. Most LLMs are not able to find alternative paths along simple graphs, indicating that their misleading explanations aren’t produced by only logical deductions using complex knowledge. These findings shed light on the limitations of the black-box explanation setting and allow us to provide advice on the safe usage of LLMs.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong explaining capabilities. They can generate texts containing logical steps towards solving complex problems (Trinh et al., 2024; Sprague et al., 2023), and incorporate users’ situations (Mondal et al., 2024; Zhu et al., 2023; Zhou et al., 2024). The literature in NLP has many works showing LLMs can generate explanations that are convincing (Wiegreffe et al., 2022), informative (Chen et al., 2023a) and faithful (Lyu et al., 2023) to levels that are comparable to human-written explanations. These abilities lead to a promising use case: LLMs as assistants that explain problems to us.
Sounds like we have a convenient use case: a user can pass a complex problem and its answer to an LLM, together with a suitably-formatted prompt. Hopefully, this “black-box” approach will let the LLM generate an explanation that is instructive and helpful for us to understand the problem. Hopefully, this use case would facilitate education, understanding, and decision-making in the ocean of complex data. However, as we will show in this paper, this approach may backfire. Specifically, the explanations generated by this “black-box” approach may encourage humans to believe in the incorrect explanandum. We refer to this problem adversarial helpfulness.
We first try to gauge the extent of this adversarial helpfulness problem. We let human annotators rate the extent they are convinced of an incorrect answer, before and after seeing an LLM-generated explanation. On a commonsense reasoning task, their ratings are significantly improved by explanations generated by three commercial-tier LLMs. We also use three top-ranking LLMs as proxy evaluators to repeat the survey, and observe similar effects.
@@
Figure 1: Examples of LLM-generated explanations towards incorrect labels in a commonsense QA question (above) and an NLI question (below). We mark the persuasion strategies with color-coded angle brackets. To save space, the parts that do not contain persuasion strategies are omitted with […].
Then, we proceed with studying the strategies of the adversarially helpful explanations. We identify ten strategies that we consider relevant to the explanations being adversarially helpful, and examine the strategies used by LLMs in generating these explanations. We find alarmingly high frequencies. For example, over 90% of the explanations in inference problems involve reframing to varying extents, and over 60% of the explanations in the commonsense problems involve the selective presentation of either the facts or the evidence.
Then we analyze the explanations from a reason graph perspective. Are these LLMs able to generate these adversarially helpful explanations because they are able to navigate through complex knowledge, like finding an alternate path in a graph? We set up a symbolic reasoning task, where we ask the LLMs to figure out an alternate path that leads to a specified destination node. This is an abstraction for “explaining an incorrect answer”. We find that the weaker models are unable to complete this task, especially when the graph complexity increases. These indicate that the generation of adversarially helpfulness explanations may involve more than the abilities in deductive reasoning and logical inference, which matches the prior observation that additional strategies (e.g., reframing) are used.
The findings shed light on the limitations of the black-box explanation setting, and motivate us to provide recommendations for safer usage of LLMs as explaining assistants. The access to all analysis scripts and data is open at https://github.com/ZiningZhu/adversarial_helpfulness.
@@ 2 Related Works
Recently, a number of works have leveraged the reasoning abilities of LLMs: chain-of-thought (Nye et al., 2021; Wei et al., 2022), tree-of-thought (Yao et al., 2024), graph-of-thought (Besta et al., 2023) and everything-of-thought (Ding et al., 2023) are representative examples. We follow their approaches to elicit the reasoning abilities from LLMs.
@@
Model-generated explanations can have significant impacts on both human users, for example in mitigating misinformation (Hsu et al., 2023; Si et al., 2024), rescaling human judgments (Wadhwa et al., 2023) and understanding models’ behavior (Hase & Bansal, 2020; Chen et al., 2022). Researchers have tried to use explanations for model development, yet there are mixed results (Saha et al., 2024; Im et al., 2023).
@@
The explainability research of AI has identified pitfalls (those unintended effects) versus dark patterns (those resulting from intended misuse) (Ehsan & Riedl, 2021). The adversarial helpfulness of LLM-generated explanation is an example of explainability pitfalls. The explanations towards models have been known to lead to problems including over-trust (Jacovi et al., 2021), over-reliance (Chen et al., 2023b) and incorrect calibrations (Zhang et al., 2020). Specific to the explanations in natural language, these explanations are selective and can be subjective or misleading (Kunz & Kuhlmann, 2024; Xu et al., 2023b) or unreliable (Ye & Durrett, 2022). Adversarial helpfulness is a different problem, as we’ll elaborate in the discussion section. LLM-generated natural language explanations about commonsense problems are perceived to have comparable attributes (generality, factuality, grammatical correctness, informativeness, acceptability) regardless of the correctness of the problems (Wiegreffe et al., 2022); we focus on the cases where the incorrect problems are explained in a post-hoc black-box manner.
@@
LLMs with self-explanations can be biased towards incorrect answers by superficial patterns like the ordering of choices (Turpin et al., 2024). Our paper considers a different setting: instead of planting superficial patterns in the demonstrations, we instruct the LLMs to explain an incorrect answer in a zero-shot manner, resembling how a user would use the LLM as an explaining assistant. This paper is also relevant to the jailbreaking/red-teaming of LLMs (Zou et al., 2023; Zeng et al., 2024; Deng et al., 2023; Ganguli et al., 2022). Instead of developing the jailbreaking or defending algorithms, we focus on the problem itself, and study the strategies adopted by the LLMs when generating adversarially helpful explanations.
@@ 3 Experiment setup
Two datasets are used: ECQA (Aggarwal et al., 2021a) and SNLI (Bowman et al., 2015). For ECQA, we sampled 500 problems and hand-pick a “second-best answer” that we consider to be only slightly worse than the correct answer designated by the dataset. Anecdotally, we found that those more abstract problems lead to more significant “adversarial helpfulness” explanations, so we skipped some concrete problems. For SNLI, we sampled 300 problems from each of the datasets with label Entailment and Contradictory, respectively. We found it was almost impossible to write sufficiently logical arguements (with humans or LLMs) that an Entailment sentence pair should be Contradictory (or vice versa), so we only considered the cases that argue for a Neutral label.
@@
We consider the following models as evaluators: Vicuna-33B-v1.3 (Chiang et al., 2023), WizardLM-70B-V1.0 (Xu et al., 2023a), Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). These are the top-ranking [^1] As of February 2024. Note that Qwen1.5-72B-Chat (Bai et al., 2023) and Yi-34B-Chat (AI et al., 2024) are also ranked on the top places, but we only included one LLM per model size to reduce the computation cost. models on the lmsys Chatbot Arena leaderboard (Zheng et al., 2023) that provide weight access. The weight access is important, since it allows us to query the probabilities in setting up the automatic evaluation — we will describe the setup in Section 5.
@@
The “adversarially helpful” explanations are generated by prompting the LLM explainer to explain an incorrect answer to a question. Figure 2 lists the prompting templates we used. In our preliminary experiments, we found other prompting templates can lead to explanations demonstrating a similar “adversarially helpful” behavior, as long as we specifically instruct the model to convince the audience of the incorrect answer.
@@ 5 Automatic evaluation of adversarial helpfulness
Note that the utility of LLM that evaluates the convincingness should be taken with caution. First, LLMs have demonstrated evidence to model humans’ thoughts (i.e., “theory-of-mind” modeling) (Kosinski, 2024), but the actual capability is being debated. Second, in several cases, the LLM-given ratings show “degeneration” trends. For example, in NLI $E\rightarrow N$, all C_after scores computed by Mixtral-8x7B are the same, resulting in an average score of 3.00. This indicates that the scores given by the proxy evaluators might be affected by the dataset artifacts, in addition to the contents.
@@
When getting down to the item-wise level, the evaluators show varying trends. First, humans correlate weakly with the proxy models. On the dataset where human annotator results are available, we compute the Pearson correlations between the averaged human results and the evaluator models. None of the correlations are significant, indicating that the evaluator models show very different fluency, correctness and convincingness ratings from the human annotators. Second, human results show poor inter-annotator agreements. This is because the MTurk platform distributes the annotation tasks to more than three annotators. Third, the proxy models assign different scores. We compute the Cohen’s Kappa between each pair of the proxy models. None of them have a value larger than 1 for any score, indicating poor agreement between the evaluators. This is because the proxy models have different “baseline marking guidelines”, as is illustrated by the drastically different mean scores in Table 1.
@@ 6 Strategies toward adversarial helpfulness
The recent literature involves many taxonomies of persuasion strategies. For example, Dimitrov et al. (2021) identified strategies in social media texts. Piskorski et al. (2023) considered news. Zeng et al. (2024) considered 40 techniques for jailbreaking LLMs as well as other NLP tasks. Inspired by these works, we identify ten persuasion strategies that we consider particularly relevant to LLM-generated explanations. A brief summarization of each strategy is included in Appendix B.
We use GPT-4 Turbo (gpt-4-0125-preview) to detect the persuasion strategies used in the explanations. Appendix B lists the prompt. If no strategy is detected, the json object parsed from the LLM’s response would contain zero in all ten entries. Figure 1 lists two examples of the identified strategies.
@@
Table 2 lists the frequencies of the persuasion strategies adopted by the three explainer models. We observe the following trends from the results.
First, the sheer numbers of frequency are alarming. For the commonsense questions, more than $70\%$ of the explanations highlight the comparative advantage towards the incorrect answers. For inference tasks, more than $90\%$ of the explanations involve reframing of the questions. While these results are only on the commonsense and inference datasets, we expect that the real-world applications of natural language explanations frequently include these persuasion strategies. Do LLM explanations exhibit these strategies for correct problems? We hypothesize an affirmative answer, since LLMs have these persuasion capabilities. However, we argue that a safe explainer should utilize few or none of these persuasion strategies in explanation, especially when the explanandum involves an incorrect problem.
Second, the three explainer models have common trends in applying persuasion strategies. The LLM-generated explanations appear overly confident. [^2] This may related to the overly-confident tone in our prompts, as listed in Figure 2. Strategies like selective evidence, and selective fact presentation are frequently used. [^3] Relatedly, selectivity is a desirable feature in human explanations (Lombrozo & Liquin, 2023). The strategies like appeal to authority and analogical evidence are very infrequently used in any of the models. These indicate that safeguarding against adversarially helpful explanations might benefit from the identification and prevention of a finite set of persuasion strategies.
@@
Future Works
This section presents an exploratory analysis of the explanation strategies, opening up future research directions. First, the cause of each strategy can be analyzed, by for example correlating each persuasion strategy and linguistic markers like syntactic complexity. Second, how each of the strategies affects the adversarial helpfulness can be studied in future works. Third, how each of the persuasion strategies can be mitigated, either by prompting techniques or by model alignment, would be promising future works.
@@ 10. Complex inference 4 1 4 9 8 5 3 7 3
Table 2: The frequencies (out of 100) of persuasion strategies adopted by explainer models. The three strategies with the highest frequencies per column are marked in bold font.
@@ 7 A structural analysis towards adversarial helpfulness
We consider the process of explanation to be an instance of path finding on a graph. In this path-finding problem, we find a path from the “root” node to the “answer” node. The explanation serves as the path that connects the known (the “root” node) to the unknown (the “answer” node).
@@
LLMs have limited capabilities in capturing the structural information in graphs (Huang et al., 2023a). Based on the intuition, we attach two simplification assumptions. First, all branches have the same path length. Second, each non-root node uniquely appears on only one path. Figure 4 shows an example of such graph with $B=3$ branches of $L=4$ path.
@@ 8 Discussion
There are intricate differences between these two improvement areas of LLM-generated explanations. Adversarial helpfulness refers to the explainer’s behavior that rationalizes an incorrect problem. (Un)faithfulness, on the other hand, is not tied to the correctness of the problem to be explained. For example, an explanation for a decision-making model can be unfaithful if the stated reasons differ from the reasoning process of this model.
@@
Plausibility is different from faithfulness — for an extended discussion of their difference please refer to Agarwal et al. (2024). Here we consider that adversarial helpfulness overlaps with plausibility; the latter indicates a potential negative impact on humans, whereas the former does not. Plausibility is a feature, but adversarial helpfulness is a bug.
Adversarial helpfulness vs hallucination
Hallucination refers to the undesirable phenomena of natural language generation (NLG) systems generating unfaithful or nonsensical texts (Ji et al., 2023). Adversarial helpfulness describes phenomena from a much smaller scope: those where the explanations facilitate the belief of the incorrect explanandum.
@@
We provide three recommendations here.
First, let humans make decisions, not AIs. We can use the AI explainer as a “Prudence” model (Miller, 2023). A “Prudence” model provides evidence supporting human’s decisions, without directly giving us an answer. The alternative, “Bluster” model, recommends an answer, optionally with the rationales supporting that answer. Since the rationales supporting the answer can be adversarially helpful, it is important to delegate the least possible amount of “decision-making” responsibility to the explainers (especially those based on generative AI). [^5] Also note that we should avoid overloading the human users with information about the model, since exposing too much model-related details could make humans less able to detect models’ mistakes (Poursabzi-Sangdeh et al., 2021).